# Desynchronization Index: A New Connectivity Approach for Exploring Epileptogenic Networks

### v2

@@ Abstract
Stereoelectroencephalography (SEEG) is an invasive surgical procedure to record the electrical activities in cortical brain regions, aiming at identifying the Epileptogenic Zone (EZ) in patients with drug-resistant epilepsy. To improve the accuracy of the EZ definition, SEEG analysis can be supported by computational tools, among which the Epileptogenic Index (EI) represents the most common solution. However, the scientific community has still not found an agreement on which quantitative biomarkers fully characterize the cortical sites within the EZ. In this work, we design a new algorithm, named Desynchronization Index (DI), to assist neurophysiologists in SEEG interpretation. Our framework hypothesizes that the EZ is identified by the cortical sites that get abnormally desynchronized from the network during the seizure onset. We test the proposed method over a SEEG dataset of 11 patients, comparing its accuracy in terms of EZ localization against the EI algorithm and clinical ground truth. Our results indicate that the DI algorithm underscores specific connectivity dynamics that can hardly be identified with a pure visual analysis, increasing system sensitivity in EZ localization with respect traditional methods.
Index Terms:
@@
This paper has been submitted to IEEE Transactions on Neural Systems and Rehabilitation Engineering
I Introduction
@@
Nowadays, SEEG interpretation is mainly based on visual analysis, focusing on the early phase of the seizure discharge. The final goal is to define an EZ that can be focal or distributed among multiple cortical structures, a scenario falling under the concept of epileptogenic networks [5]. The above process is particularly complex since SEEG signals include hundreds of components, whose analysis requires the involvement of highly specialized neurophysiologists. For this reason, in the last years, SEEG interpretation has started being supported by computational tools apt at characterizing the EZ and, more in general, the dynamics leading to seizure generation [6]. The most common methods analyze the SEEG power spectrum, searching for the cortical sites generating fast oscillations, i.e., electrical activities with frequencies in the gamma range ($[30,100]$ Hz), as done in the cases of Epileptogenic Index (EI) [7] and Epileptogenic Maps (EMs) [8].
@@
Although multiple quantitative approaches have been developed for analyzing the connectivity dynamics of epileptogenic networks, the literature often presents discordant findings. Most works agree that the ictal phase (i.e., the period involving the seizure discharge) is preceded by a connectivity reduction, followed by the opposite phenomenon during the seizure propagation [10]. Recently, it has been hypothesized that the EZ presents an abnormal inward information flow during the inter-ictal phase [11], and the epileptogenic level of each cortical site can be assessed by the connection exerted during the epileptic discharge [12].
In this work, we present a new framework for supporting neurophysiologists in SEEG interpretation. Specifically, we develop a connectivity-based algorithm, named Desynchronization Index (DI), that exploits the Phase Transfer Entropy (PTE) to estimate the relations between brain structures and detect connectivity anomalies in the resulting time-varying network. According to our approach, the EZ is identified by the cortical sites that present an independent behavior during the ictal phase. Overall, the main contributions of this manuscript consist of the following points:
@@ II Related Work
Alongside designing the EI, Bartolomei and colleagues analyze SEEG data by the nonlinear regression coefficient (h2) [24], and show that the epileptogenic network is associated with reduced connectivity in the period that foresees the ictal phase  [25]. While the h2 coefficient is agnostic to the signal frequencies, most connectivity models discern cortical relationships depending on the frequency bands. Following this principle, the authors of [26] estimate connectivity via linear correlation and observe that seizure frequency is proportional to the EZ outward connectivity. Similar methodologies are given in [27], whose authors combine pairwise correlation with Euclidean distance to estimate the EZ extension, and [28], where graph theory is exploited to identify the network hubs most contributing to the seizure propagation.
@@
Recently, the use of Spearman rank correlation [29] to assess the coupling between brain structures has shown good results for both the detection of the EZ and the prediction of surgical outcomes [30]. Other connectivity-based approaches correlate the phase of slow oscillations ($<30$ Hz) with the amplitude of fast oscillations ($>30$ Hz), as occurred for the Phase Slope Index (PSI) and the Phase Amplitude Coupling (PAC) [31]. Exploiting the PSI, [32] shows that the EZ presents higher outward connectivity during the seizure onset than the resting state. Similar results are observed in [33], where the PAC between high (gamma) and low (delta and theta) frequency bands is computed. The authors of [34] seem to contradict these findings and, using the Partial Direct Coherence (PDC) [35] as a connectivity measure, assess that the EZ presents higher inward connectivity during seizure generation.
@@
In the last few years, the popularity of Machine Learning (ML) has encouraged many researchers to apply data-driven techniques to epileptogenic networks. For instance, the authors of [36] exploit hypergraph learning for automatically detect SEEG fast oscillations, while [37] investigates different patient-specific ML models for analyzing scalp electroencephalography signals. Besides, [38] exploit a Support Vector Machine (SVM) model to integrate multiple electrophysiological features, including both fast oscillations and low frequency suppression, for enhancing EZ detection. This supports the fact that ictal channels are not identified by specific frequency components, and more holistic methodologies are required to describe seizure generation. On the other hand, [38] and similar works consider as ground truth the area resected by epilepsy surgery, which is usually far larger than the EZ to maximize the probability of seizure reduction. In general, the high complexity of SEEG analysis makes it is extremely hard to build supervised datasets, which prevents ML solutions from obtaining reliable performance in this field [39].
@@
Although the large variety of works implementing connectivity-based metrics for SEEG interpretation, results from the literature do not lead to an agreement on which biomarkers characterize the EZ. Initial works show that the ictal transition is characterized by a connectivity reduction, but the EZ is often identified by the cortical sites with a higher coupling during the seizure propagation. In this work, we try to overcome such questions, by quantifying the epileptogenicity level of SEEG sites in terms of desynchronization. With this latter, we denote the tendency of a cortical site to assume an independent electrical behavior that cannot be inhibited by neighboring brain structures.
III Analytical Model
@@ III-A Epileptogenic Index
The above formulation stands out from the original EI algorithm by reducing the number of parameters to be manually set. In particular, $t_{\text{start}}$ and $t_{\text{end}}$ derive from the clinical interpretation of the SEEG recording and denote the period during which the epileptic discharge takes shapes. Instead, $t_{\text{base}}$ should be defined in such a way to ensure that the statistics $\mu_{E_{x}}$ and $\sigma_{E_{x}}$ are as accurate as possible. Hence, the only parameters to be tuned are the maximum number $M$ of epileptogenic channels, the interval $\delta$ over which the tonicity is computed (used in (4)), and the weight $\gamma$ of the CUSUM control chart (used in (2)), which trades off between the false alarm and miss-detection probabilities.
III-B Effective Connectivity Model
@@
The proposed methodology quantizes the phase values to reduce the influence of volume conduction on the connectivity model. A similar same principle is adopted by the PLI algorithm, which, however, does not consider the effectiveness of the network connection [18]. In addition, PLI presents critical pitfalls, including underestimating those connections associated with small propagation delays or with signals whose phase difference distribution fluctuates around zero. Our technique tries to act as a compromise between PLI and PLV, which, notably, suffers the effect of volume conduction and leads to many false connectivity phenomena.
@@ III-C Desynchronization Index
where $\alpha\in[0,1]$ is a tuning parameter that determines the rate at which older information influences the EWMA statistic. In particular, as $\alpha\rightarrow 1$, the function gives more importance to the most recent observations, better capturing rapid process changes at the cost of an increased false alarm probability.
In the following, we write $\mathcal{T}(t)$ to indicate the distribution of the inter-channel connections $T_{x\rightarrow y}(t)$, $\forall\,x,y\in\mathcal{N}$, at time $t$. Hence, we denote by $E[\mathcal{T}(t)]$, $P_{75}[\mathcal{T}(t)]$, and $P_{25}[\mathcal{T}(t)]$ the expectation, the $75$th percentile, and the $25$th percentile of $\mathcal{T}(t)$. We use the EWMA algorithm, as defined in Eq.(10), to compute a smoothed version of the above statistics as well as of each connection $T_{x\rightarrow y}(t)$, $\forall\,x,y\in\mathcal{N}$. Then, we define $\mathcal{N}_{x}^{\text{in}}(t)$ as the set of channels presenting abnormal inward connectivity with respect $x$ at time $t$:
@@
In other words, we consider abnormal those connections that go outside the inter-quartile range of the time-varying network distribution. This, on the one hand, allows us to implicitly customize the framework to each patient without the need to define additional parameter as in [41]. On the other hand, choosing other percentile ranges for discerning abnormal connections could lead to different trade-offs in terms of accuracy, and, in the future, further investigation of how to tune such a parameter is required.
TABLE I: Demographic and clinical data.
@@ 1 Male 39 33 Weekly Right temporal Temporal pole Lobectomy Seizure free HS 1
2 Male 24 1 Daily Left orbito-frontal Fronto-insular RF-TC 1 seizure in 1 year Not available
@@ 6 Male 48 22 Weekly Right temporal Orbito-frontal, and insular Lobectomy Seizure free HS 1
7 Male 36 9 Weekly Right temporo-occipital Temporo-basal and mesial RF-TC Not available Not available
8 Female 49 12 Monthly Left temporal Temporo-lateral Lobectomy Not available HS 1
@@ 11 Male 24 18 Weekly Right temporal Temporal pole RF-TC Seizure free Tumor
We define the actual $\psi_{x}^{\text{in}}(t)$ and expected $\hat{\psi}_{x}^{\text{in}}(t)$ inward connection density of channel $x$ at time $t$ as follow:
$\psi_{x}^{\text{in}}(t)=\sum_{y\in\mathcal{N}_{x}^{\text{in}}(t)}\mathcal{E}(T_{y\rightarrow x},t);$ (12)
$\hat{\psi}_{x}^{\text{in}}(t)=|\mathcal{N}_{x}^{\text{in}}(t)|\cdot\mathcal{E}(E[\mathcal{T}],t).$ (13)
We can now define the desynchronization level of a channel $x\in\mathcal{N}$ at time $t$ as
$D_{x}(t)=\sqrt{\hat{\psi}_{x}^{\text{in}}(t)}-\sqrt{\psi_{x}^{\text{in}}(t)}.$ (14)
In particular, at each time $t$, $D_{x}(t)$ represents the tendency for channel $x\in\mathcal{N}$ to assume an independent behavior.
We apply the CUSUM chart to the desynchronization level, thus obtaining $\Gamma(D_{x},t)$ for each channel $x\in\mathcal{N}$ and for each time $t\in[t_{\text{start}},t_{\text{end}}]$. Hence, using equations (3) and (4), we can compute the activation times $t_{D_{x}}$ and the tonicities $c_{D_{x}}$, $\forall$ $x\in\mathcal{N}$, considering the desynchronization level instead of the energy ratio. Following the same approach presented in III-A, we define the set $\mathcal{N}_{D}$, including the $M$ channels showing the strongest changes in terms of connectivity. Finally, for each channel in $\mathcal{N}_{D}$, we define the DI values as
@@
Appreciably, the DI algorithm follows the same settings of EI, making it necessary to tune only three parameters, namely the maximum number $M$ of epileptogenic channels, the interval $\delta$, and the weight $\gamma$.
IV Evaluation Methodology
In this section, we present the clinical dataset where our tool is evaluated, describing the SEEG recording process and the patient characteristics. Then, we showcase how the EI and DI algorithms are implemented and describe how we evaluate their performance in the detection of the EZ.
IV-A Clinical Dataset
This study considers a cohort of $12$ consecutive patients that were monitored through SEEG at IRCCS Institute of Neurological Sciences of Bologna from January 2022 to June 2024. The protocol for the study was approved by the local ethics committee (protocol number 741-2021, committee code 97338), and written informed consent was obtained from each patient. In all the cases, the SEEG implant included multiple electrodes, each presenting 555518181818 recording sites, named contacts; the number and location of the electrodes were patient-tailored, while each contact was $22$ mm in length, and separated by $1.5$ m from neighboring contacts.
@@ IV-A Clinical Dataset
From the initial dataset, a single patient was excluded since the SEEG recording did not allowed the localization of the EZ and no epilepsy surgery was conducted. The final cohort comprises 7 males and 4 females with an average age of 36.6 years (range 24-49) at the recording time and 8.8 years (range 1-22) at the time of the disease onset. The clinical data pointed to a probable EZ in the temporal lobe in 7 cases (5 right, 2 left), in the left-frontal lobe in a single case, and in temporo-occipital lobe in the remaining cases (2 right, 1 left). At the time of SEEG implantation, seizure frequency ranged from daily to monthly, depending on the patient. Radio-frequency thermocoagulations enabled the control of seizures in a single patient, while the remaining underwent anterior-temporal lobectomy. Histopathological analysis led to clinical results compatible with hippocampal sclerosis in 4, and cortical dysplasia in 1 cases. Demographic and clinical details of the studied population are reported in Tab. I.
HS stays for hippocampal sclerosis FCD for focal cortical dysplasia, RF-TC for radio-frequency thermocoagulation, FOI for fronto-orbito-insular.
IV-B Performance Evaluation
For each patient, we consider the channels belonging to the EZ as the ground truth. The selection of the EZ channels was performed by two board-certified neurophysiologists (L.D-V. and E.P.), who were blinded to the output of our computational framework. A single seizure per patient was selected, and an SEEG epoch of $T_{\text{epoch}}=200$ seconds is analyzed for each seizure, considering the period before the complete diffusion of the epileptic discharge. If multiple seizures were available, we chose to discard the seizures recorded during the first two days of the monitoring period (during which the pharmacological treatment is still active) and those occurring during a seizure cluster.
Before running the EI and DI algorithms, the SEEG signals are pre-processed through a notch filter centered at $f_{\text{notch}}=50$ Hz. In the analysis, we exclude the sites exploring the White Matter (WM) since they are not sources of electric signals. The selection of the WM channels was performed by a board-certified neurophysiologist (L.F.), according to the electrical activity of the cortical sites and their localization in the SEEG multimodal scene. No other cortical sites are excluded from the analysis, making our framework agnostic to the specific patient and seizure to be analyzed.
The output of the EI and DI values are normalized in the $[0,1]$ interval, and a threshold $\eta$ (in the same range) is set for classifying a channel as epileptogenic. Hence, the following performance metrics are considered:

### v3

@@ Abstract
Objective: In this work, we propose a new computational framework to assist neurophysiologists in Stereoelectroencephalography (SEEG) analysis, with the final aim of improving the definition of the Epileptogenic Zone (EZ) in patients with drug-resistant epilepsy. Method: We consider the Phase Transfer Entropy (PTE) model to estimate the effective connectivity between SEEG channels. Hence, we design an algorithm, named the Desynchronization Index (DI), that classifies as epileptogenic the channels that show independent behavior with respect to the rest of the network during the seconds preceding the seizure propagation. Results: We test the proposed DI algorithm against Epileptogenic Index (EI) on a clinical dataset of 11 patients, considering the neurophysiological evaluation of the EZ as the clinical ground truth. Our results denote that DI overcomes EI in terms of area under the ROC curve (AUC=0.80 vs AUC=0.74), while combining the two algorithms as a unique tool leads to the best performance (AUC=0.87). Conclusions: The DI algorithm underscores connectivity dynamics that can hardly be identified with a pure visual analysis, increasing the accuracy in the EZ definition compared to traditional methods. Significance: The integration of connectivity- and energy-based metrics, as the one we propose, can lead to the definition of a new effective biomarker of the EZ, reducing the burden required by the SEEG review in the case of extensive implants and improving our understanding of the dynamics leading to the generation of seizures.
{IEEEkeywords}
@@
This paper has been submitted to IEEE Transactions on Biomedical Engineering
1 Introduction
[…]
@@ 1 Introduction
Nowadays, SEEG interpretation is mainly based on visual analysis, focusing on the early phase of the seizure discharge. The final goal is to define an EZ that can be focal or distributed among multiple cortical structures, a scenario falling under the concept of epileptogenic networks [5]. The above process is particularly complex since SEEG signals include hundreds of components, whose analysis requires the involvement of highly specialized neurophysiologists. For this reason, in the last years, SEEG interpretation has started being supported by computational tools apt at characterizing the EZ and, more in general, the dynamics leading to seizure generation [6]. The most common methods analyze the SEEG power spectrum, searching for the cortical sites generating fast oscillations [7], i.e., electrical activities with frequencies in the gamma range ($[30,100]$ Hz), as done in the cases of Epileptogenic Index (EI) [8] and Epileptogenic Maps (EMs) [9].
@@
Although multiple quantitative approaches have been developed for the study of epileptogenic networks, the literature often presents discordant findings. Most works agree that the ictal phase (i.e., the period involving the seizure discharge) is preceded by a connectivity reduction, followed by the opposite phenomenon during the seizure propagation [11]. Recently, it has been hypothesized that the EZ presents an abnormal inward information flow during the inter-ictal phase [12], and the epileptogenic level of each cortical site can be assessed by the connection exerted during the epileptic discharge [13].
In this work, we present a new framework for supporting neurophysiologists in SEEG interpretation. Specifically, we design a connectivity-based algorithm, named Desynchronization Index (DI), that exploits the Phase Transfer Entropy (PTE) to estimate the relations between brain structures and detect connectivity anomalies in the resulting time-varying network. According to our approach, the EZ is identified by the cortical sites that present an independent behavior during the ictal phase. Overall, the main contributions of this manuscript consist of the following points:
@@ 2 Related Work
In addition to designing EI, Bartolomei et al. analyze the SEEG data using the nonlinear regression coefficient (h2) [25], and show that the epileptogenic network is associated with reduced connectivity in the period that anticipates the ictal phase  [26]. While the h2 coefficient is agnostic to the signal frequencies, most connectivity models discern cortical relationships depending on the frequency bands. Following this principle, the authors of [27] estimate connectivity via linear correlation and observe that seizure frequency is proportional to the EZ outward connectivity. Similar methodologies are given in [28], whose pairwise correlation is combined with Euclidean distance to estimate the EZ extension, and [29], where graph theory is exploited to identify the network hubs that most contribute to the seizure propagation.
@@
Recently, the use of Spearman rank correlation [30] to assess the coupling between brain structures has shown good results for both the detection of the EZ and the prediction of surgical outcomes [31]. Other connectivity-based approaches associate the phase of slow oscillations ($<30$ Hz) with the amplitude of fast oscillations ($>30$ Hz), as occurred for the Phase Slope Index (PSI) and the Phase Amplitude Coupling (PAC) [32]. Exploiting the PSI, [33] shows that the EZ presents higher outward connectivity during the seizure onset than the resting state. Similar results are observed in [34], where the PAC between high (gamma) and low (delta and theta) frequency bands is computed. The authors of [35] seem to contradict these findings and, using the Partial Direct Coherence (PDC) [36] as a connectivity measure, assess that the EZ presents higher inward connectivity during seizure generation.
@@
In recent years, the popularity of Machine Learning (ML) has encouraged many researchers to apply data-driven techniques to epileptogenic networks. For example, the authors of [37] exploit hypergraph learning to automatically detect SEEG fast oscillations, while [38] investigates different patient-specific ML models for analyzing scalp electroencephalography signals. In addition, [39] exploit a Support Vector Machine (SVM) model to integrate multiple electrophysiological characteristics, including both fast oscillations and low-frequency suppression, to improve accuracy in EZ detection. This supports the fact that ictal channels are not identified by specific frequency components, and more holistic methodologies are required to describe the generation of seizures. On the other hand, [39] and similar works consider as ground truth the area resected by epilepsy surgery, which is usually far larger than the EZ to maximize the probability of seizure reduction. In general, the high complexity behind SEEG analysis makes it extremely difficult to build supervised datasets, which prevents ML solutions from obtaining reliable performance in this field [40].
@@
Although the large variety of works implementing connectivity-based metrics for the interpretation of SEEG, the results of the literature do not lead to an agreement on which biomarkers characterize EZ. Initial works show that the ictal transition is characterized by a reduction in connectivity, but the EZ is often identified by cortical sites with a greater coupling during seizure propagation. In this work, we try to answer these questions by quantifying the level of epileptogenicity of SEEG sites in terms of desynchronization. With this latter, we denote the tendency of a cortical site to show independent electrical behavior that cannot be inhibited by neighboring brain structures.
3 Analytical Model
@@ 3.1 Epileptogenic Index
The above formulation stands out from the original EI algorithm by reducing the number of parameters to be manually set. In particular, $t_{\text{start}}$ and $t_{\text{end}}$ denote the period during which the epileptic discharge takes shapes. Instead, $t_{\text{base}}$ should be defined in such a way to ensure that the statistics $\mu_{E_{x}}$ and $\sigma_{E_{x}}$ are as accurate as possible. Hence, the only parameters to be tuned are the maximum number $M$ of epileptogenic channels, the interval $\delta$ over which the tonicity is computed, and the weight $\gamma$ of the CUSUM control chart, which trades off between the false alarm and miss-detection probabilities.
3.2 Effective Connectivity Model
@@
The proposed methodology quantizes the phase values to reduce the influence of volume conduction on the connectivity model. A similar same principle is adopted by the PLI algorithm, which, however, does not consider the effectiveness of the network connection [19]. In addition, PLI presents critical pitfalls, including underestimating those connections associated with small propagation delays or with signals whose phase difference distribution fluctuates around zero. Hence, PTE tries to act as a compromise between PLI and PLV, which, notably, suffers the effect of volume conduction and leads to many false connectivity phenomena.
@@ 3.3 Desynchronization Index
where $\alpha\in[0,1]$ is a tuning parameter that determines the rate at which older information influences the EWMA statistic. In particular, as $\alpha\rightarrow 1$, the function gives more importance to the most recent observations, better capturing rapid changes in the process at the cost of an increased false alarm probability.
In the following, we write $\mathcal{T}(t)$ to indicate the distribution of the inter-channel connections $T_{x\rightarrow y}(t)$, $\forall\,x,y\in\mathcal{N}$, at time $t$. Hence, we denote by $M[\mathcal{T}(t)]$, $P_{75}[\mathcal{T}(t)]$, and $P_{25}[\mathcal{T}(t)]$ the median, the $75$th percentile, and the $25$th percentile of $\mathcal{T}(t)$. We use the EWMA algorithm, as defined in Eq.(10), to compute a smoothed version of the above statistics as well as of each connection $T_{x\rightarrow y}(t)$, $\forall\,x,y\in\mathcal{N}$. Hence, we define $\mathcal{N}_{x}^{\text{in}}(t)$ and $\mathcal{N}_{x}^{\text{out}}(t)$ as the sets of channels presenting abnormal connectivity toward and from channel $x$ at time $t$:
@@
In other words, we consider abnormal those connections that go outside the inter-quartile range of the time-varying network distribution. This, on the one hand, allows us to implicitly customize the framework to each patient without the need to define additional parameters as in [42]. On the other hand, choosing other percentile ranges for discerning abnormal connections could lead to different trade-offs in terms of accuracy, and, in the future, further investigation of how to tune such a parameter is required.
Table 1: Demographic and clinical data.
@@ 1 Male 39 33 Weekly Right temporal Temporal pole Lobectomy Seizure free HS 1
2 Male 24 1 Daily Left orbito-frontal Fronto-insular RF-TC Seizure reduction Not available
@@ 6 Male 48 22 Weekly Right temporal Orbito-frontal, and insular Lobectomy Seizure free HS 1
7 Male 36 9 Weekly Right temporo-occipital Temporo-basal and mesial RF-TC Seizure reduction Not available
8 Female 49 12 Monthly Left temporal Temporo-lateral Lobectomy Seizure free HS 1
@@ 10 Female 38 13 Monthly Right temporo-occipital Temporo-mesial RF-TC Not available Not available
We define the actual $\psi_{x}^{\text{in}}(t)$ and expected $\hat{\psi}_{x}^{\text{in}}(t)$ inward connection density of channel $x$ at time $t$ as
$\psi_{x}^{\text{in}}(t)=\sum_{y\in\mathcal{N}_{x}^{\text{in}}(t)}\mathcal{E}(T_{y\rightarrow x},t),$ (12)
$\hat{\psi}_{x}^{\text{in}}(t)=|\mathcal{N}_{x}^{\text{in}}(t)|\cdot\mathcal{E}(M[\mathcal{T}],t).$ (13)
In a specular way, we define $\psi_{x}^{\text{out}}(t)$ and $\hat{\psi}_{x}^{\text{out}}(t)$ as
$\psi_{x}^{\text{out}}(t)=\sum_{y\in\mathcal{N}_{x}^{\text{out}}(t)}\mathcal{E}(T_{x\rightarrow y},t),$ (14)
$\hat{\psi}_{x}^{\text{out}}(t)=|\mathcal{N}_{x}^{\text{out}}(t)|\cdot\mathcal{E}(M[\mathcal{T}],t).$ (15)
We now define the desynchronization level of channel $x\in\mathcal{N}$ at time $t$ as
$D_{x}(t)=\sqrt{\hat{\psi}_{x}^{\text{in}}(t)}-\sqrt{\psi_{x}^{\text{in}}(t)}+\sqrt{\hat{\psi}_{x}^{\text{out}}(t)}-\sqrt{\psi_{x}^{\text{out}}(t)},$ (16)
where the square root operator is used to stabilize the variance of the process [42]. The first part of the above equation depends on the inward connection of $x$, while the second refers to the influence that $x$ exerts on other channels. At each time $t$, $D_{x}(t)$ represents the tendency of $x$ to self-isolate from the rest of the network.
We apply the CUSUM chart to the desynchronization level, thus obtaining $\Gamma(D_{x},t)$ for each channel $x\in\mathcal{N}$ and for each time $t\in[t_{\text{start}},t_{\text{end}}]$. Hence, using equations (3) and (4), we can compute the activation times $t_{D_{x}}$ and the tonicities $c_{D_{x}}$, $\forall$ $x\in\mathcal{N}$, considering the desynchronization level instead of the energy ratio. Following the same approach presented in 3.1, we define the set $\mathcal{N}_{D}$, including the $M$ channels showing the strongest changes in terms of connectivity. Finally, for each channel in $\mathcal{N}_{D}$, we define the DI values as
@@
The DI algorithm follows the same settings of EI, making it necessary to tune only the maximum number $M$ of epileptogenic channels, the interval $\delta$, and the weight $\gamma$.
4 Evaluation Methodology
In this section, we present the clinical dataset in which our tool is evaluated, describing the SEEG recording process and the patient characteristics. Then, we showcase how the EI and DI algorithms are implemented and describe how we evaluate their performance in the detection of the EZ.
4.1 Clinical Dataset
This study considers a cohort of $12$ consecutive patients that were monitored through SEEG at the IRCCS Institute of Neurological Sciences of Bologna from January 2022 to June 2024. The study protocol was approved by the local ethics committee (protocol number 741-2021, committee code 97338), and written informed consent was obtained from each patient. In all cases, the SEEG implant included multiple electrodes, each presenting 555518181818 recording sites, named contacts; the number and location of the electrodes were patient-tailored, while each contact was $22$ mm in length, and separated by $1.5$ m from neighboring contacts.
@@ 4.1 Clinical Dataset
From the initial dataset, a single patient was excluded as the SEEG record did not allow the localization of the EZ and no epilepsy surgery was performed. The final cohort comprises 7 males and 4 females with an average age of 36.6 years (range 24-49) at the recording time, and 8.8 years (range 1-22) at the disease onset. Clinical data indicated a probable EZ in the temporal lobe in 7 cases (5 right, 2 left), in the left frontal lobe in a single case, and in the temporo-occipital lobe in the remaining cases (2 right, 1 left). Radiofrequency thermocoagulations allowed control of seizures in a single patient, while the remaining patients underwent anterior-temporal lobectomy. Histopathological analysis led to clinical results compatible with hippocampal sclerosis in 4, and cortical dysplasia in 1 cases. Demographic and clinical details of the studied population are reported in Tab. 1  [^1] HS stays for hippocampal sclerosis FCD for focal cortical dysplasia, RF-TC for radio-frequency thermocoagulation, FOI for fronto-orbito-insular..
4.2 Performance Evaluation
For each patient, we consider the channels belonging to the EZ as the detection target. The selection of the EZ channels was performed by two board-certified neurophysiologists (L.D-V. and E.P.), who were blinded to the output of our computational framework. We selected a single seizure per patient and analyzed an SEEG epoch of $T_{\text{epoch}}=200$ seconds for each seizure, considering the period before the complete diffusion of epileptic discharge. If multiple seizures were available, we discarded the seizures recorded during the first two days of the monitoring period (during which the pharmacological treatment is still active) and those occurring during a cluster of seizures.
Before running the EI and DI algorithms, the SEEG signals were pre-processed through a notch filter centered at $f_{\text{notch}}=50$ Hz. In the analysis, we excluded sites that explore the White Matter (WM) since they are not sources of electric signals. The selection of the WM channels was performed by a board-certified neurophysiologist (L.F.), according to the electrical activity of the cortical sites and their location on the multimodal scene. No other cortical sites were excluded, making our framework agnostic to the specific patient and seizure to be analyzed.
The output of the EI and DI algorithms is normalized in the $[0,1]$ interval, and a threshold $\eta$ (in the same range) is set to classify a channel as epileptogenic. Hence, we consider the following performance metrics: